OpenStudy (anonymous):

can you help me simplify: 9y-3x/x-3y

OpenStudy (anonymous):

is it -3?

OpenStudy (anonymous):

=3(3y-x)/x-3y=-3* x-3y/x-3y = -3

OpenStudy (anonymous):

i followed an example earlier and it said 18y-2x/x-9y= -2

OpenStudy (anonymous):

okok! great... i see. thanks for the help @anotherxx